alternatives good fair poor
open 1 380000 70000 -400000
open 2 200000 80000 -200000
do nothing 0 0 0

Nick has plans to open some pizza restaurants, but he is not sure how many to open. He has prepared a payoff table to help analyze the situation. Nick believes there is a 40 percent chance that the market will be good, a 30 percent chance that it will be fair, and a 30 percent chance that it will be poor. A market research firm will analyze market conditions and will provide a perfect forecast (they provide a money back guarantee). What is the maximum nick should be willing to pay for perfect information? 53,000 176,000 123,000 132,000 280,000

0 0

> A

daniel johnson Thu, Dec 9, 2021 2:02 PM

Add a comment Improve this question Transcribed image text
Answer #1

Correct Answer:

$123000

Working note:

EMV for open 1 = .4*380000 + .3*70000 + .3*(-400000) = $53000

EMV for open 2 = .4*200000 + .3*80000 + .3*(-200000) = $44000

EMV for open 3 = .4*0 + .3*0 + .3*0 = 0

So, best EMV (EVWPI) = $53000

EV with perfect information (EVPI) = .4*380000 + .3*80000 + .3*0 = 176000

So,

Expected value of perfect information = EVPI - EVWPI = 176000 - 53000

Expected value of perfect information = $123000